# AI in Supply Chain Risk Assessment: A Systematic Literature Review and Bibliometric Analysis

### v4

@@ 8 Challenges and Limitations
    Data Availability and Quality: AI models in SCRA rely heavily on data. However, ensuring the availability and quality of data remains a persistent challenge. SC data often comes from various sources, which can be fragmented, incomplete, or outdated. Researchers must explore ways to consolidate, cleanse, and augment data to enhance the effectiveness of AI models.
@@
    Interpretability and Explainability: Many AI techniques, such as DL and neural networks, are often considered “black-box” models. This lack of interpretability and explainability is a significant limitation. In SCRA, stakeholders require insights into why a model makes a particular prediction or decision. Developing AI models that are not only accurate but also interpretable is crucial for their adoption in real-world SCM.
@@
    Dynamic and Evolving Risk Factors: The SC environment is dynamic and continuously evolving. New risk factors emerge, and existing ones change over time. AI models may struggle to adapt to these dynamic conditions. Researchers need to explore techniques for real-time risk assessment and adaptive models that can respond to changing risk profiles.
@@
    Integration with Existing Systems: Integrating AI solutions into existing SCM systems can be challenging. Organizations often rely on legacy systems, and ensuring compatibility and seamless integration can pose hurdles. Research in this area should focus on creating modular AI solutions that can be integrated into diverse systems with minimal disruption.
@@ 9 Future Research Directions
In light of the challenges and limitations, future research directions in AI-based SCRA are multi-faceted and have the potential to advance the state-of-the-art. The development of AI models that deliver accurate predictions and offer comprehensive transparency and interoperability should be emphasized. Techniques like Local Interpretable Model-Agnostic Explanations (LIME) and SHapley Additive exPlanations (SHAP) serve as pivotal approaches to enhance the interpretability of AI models. By employing these methods, the goal is to demystify the decision-making process of AI models, enabling stakeholders to comprehend the rationale behind specific predictions or decisions. This transparency enhances trust and facilitates more effective integration of AI-driven insights into SC risk management strategies.
@@
Combining multiple AI techniques, such as ML and expert systems, can enhance the robustness of SCRA. Future research should explore hybrid AI models that leverage the strengths of different approaches to improve prediction accuracy and reliability. Addressing data quality challenges involves multifaceted strategies. Developing robust methods for data cleansing and augmentation and ensuring high-quality data remains a critical focus. Beyond that, initiatives should be aimed at democratizing access to standardized SC data repositories. Creating and maintaining such repositories enables stakeholders across the SC spectrum to access reliable and consistent data. Increased accessibility fosters better-informed decision-making processes and promotes using standardized data sets for improved AI model training and validation. Investigating AI models capable of real-time risk assessment will be pivotal. These models should adapt to emerging risks and provide timely alerts or recommendations for risk mitigation.
@@
Research on creating integration frameworks for AI-based SCRA is essential. These frameworks should facilitate easy integration into existing SC management systems, ensuring a smooth transition to AI-driven risk assessment. Expanding the application of AI in risk assessment to various domains within the SC, including logistics, procurement, and production, offers opportunities for comprehensive risk management. Investigate AI models capable of adapting to pandemic-induced disruptions within SCs. Focus on understanding the unique dynamics and challenges posed by COVID-19, aiming to develop AI-driven strategies that enable resilience and agility in such crisis scenarios. As AI plays an increasingly significant role in decision-making, researchers should also address ethical and legal considerations. This includes issues related to data privacy, bias, and fairness in AI models. By addressing these challenges and pursuing these research directions, AI-based SCRA can make significant strides toward more effective and reliable risk management in an ever-evolving global SC landscape.
@@ 10 Conclusions
This extensive review delved into the SCRA domain and examined the complex challenges of integrating AI, especially ML. Our thorough analysis encompassed a substantial pool of 1,439 SCRA papers, ultimately narrowing our focus to a refined subset of 51 papers. This endeavor has yielded significant insights and key findings, making substantial contributions to the SCRA literature.
@@
    Our review thoroughly analyzes the existing state of research concerning the application of AI and ML techniques in SCRM. We have meticulously assessed a substantial body of literature to extract meaningful insights, specifically 51 selected articles.
@@
    Through our analysis, we have identified the AI and ML techniques that are most commonly utilized in the domain of SCRA. This clarifies the methodologies employed and underscores the trends shaping the field.
@@
    Our review highlights the key findings and trends prevalent in the literature, shedding light on the progress, developments, and major areas of focus within the realm of AI and ML in SCRM.
@@
Despite these contributions, our research has some limitations. Our data primarily originated from Google Scholar and WoS rather than Scopus, potentially introducing biases. Our focus on papers published between 2015 and 2024 presents a limitation, as it may omit earlier relevant research that could provide valuable historical context. Additionally, our restriction to English-language papers may have omitted valuable non-English research.

### v5

@@ 8 Challenges and Limitations
    Data Availability and Quality: AI models in SCRA rely heavily on data. However, ensuring the availability and quality of data remains a persistent challenge. SC data often comes from various sources, which can be fragmented, incomplete, or outdated (Schöpper and Kersten,, 2021). Researchers must explore ways to consolidate, cleanse, and augment data to enhance the effectiveness of AI models.
@@
    Interpretability and Explainability: Many AI techniques, such as DL and neural networks, are often considered “black-box” models. This lack of interpretability and explainability is a significant limitation (Singh et al.,, 2024). In SCRA, stakeholders require insights into why a model makes a particular prediction or decision. Developing AI models that are not only accurate but also interpretable is crucial for their adoption in real-world SCM.
@@
    Dynamic and Evolving Risk Factors: The SC environment is dynamic and continuously evolving (Nitsche et al.,, 2021). New risk factors emerge, and existing ones change over time. AI models may struggle to adapt to these dynamic conditions. Researchers need to explore techniques for real-time risk assessment and adaptive models that can respond to changing risk profiles.
@@
    Integration with Existing Systems: Integrating AI solutions into existing SCM systems can be challenging due to various drivers (technological, organizational, environmental, and human) and barriers (technical, organizational, economic, and human) that influence adoption (Shahzadi et al.,, 2024). Organizations often rely on legacy systems, and ensuring compatibility and seamless integration can pose hurdles. Research in this area should focus on creating modular AI solutions that can be integrated into diverse systems with minimal disruption.
@@ 9 Future Research Directions
In light of the challenges and limitations, future research directions in AI-based SCRA are multi-faceted and have the potential to advance the state-of-the-art. The development of AI models that deliver accurate predictions and offer comprehensive transparency and interoperability should be emphasized. Techniques like Local Interpretable Model-Agnostic Explanations (LIME) and SHapley Additive exPlanations (SHAP) serve as pivotal approaches to enhance the interpretability of AI models (Olan et al.,, 2024). By employing these methods, the goal is to demystify the decision-making process of AI models, enabling stakeholders to comprehend the rationale behind specific predictions or decisions. This transparency enhances trust and facilitates more effective integration of AI-driven insights into SC risk management strategies.
@@
Combining multiple AI techniques, such as ML and expert systems, can enhance the robustness of SCRA. Future research should explore hybrid AI models that leverage the strengths of different approaches to improve prediction accuracy and reliability (Mojdehi et al.,, 2025). Addressing data quality challenges involves multifaceted strategies. Utilizing larger datasets for analysis, as emphasized by (Wang et al.,, 2021), enhances the robustness of risk assessment, enabling a more comprehensive evaluation of risk factors and improving predictive accuracy. Additionally, initiatives should be aimed at democratizing access to standardized SC data repositories. Creating and maintaining such repositories enables stakeholders across the SC spectrum to access reliable and consistent data. Increased accessibility fosters better-informed decision-making processes and promotes using standardized data sets for improved AI model training and validation. Investigating AI models capable of real-time risk assessment will be pivotal. These models should adapt to emerging risks and provide timely alerts or recommendations for risk mitigation.
@@
Research on creating integration frameworks for AI-based SCRA is essential (Shahzadi et al.,, 2024). These frameworks should facilitate easy integration into existing SC management systems, ensuring a smooth transition to AI-driven risk assessment. Expanding the application of AI in risk assessment to various domains within the SC, including logistics, procurement, and production, offers opportunities for comprehensive risk management. Investigate AI models capable of adapting to pandemic-induced disruptions within SCs. Focus on understanding the unique dynamics and challenges posed by COVID-19, aiming to develop AI-driven strategies that enable resilience and agility in such crisis scenarios. As AI plays an increasingly significant role in decision-making, researchers should also address ethical and legal considerations. This includes issues related to data privacy, bias, and fairness in AI models. By addressing these challenges and pursuing these research directions, AI-based SCRA can make significant strides toward more effective and reliable risk management in an ever-evolving global SC landscape.
@@ 10 Conclusions
This extensive review delved into the SCRA domain and examined the complex challenges of integrating AI, especially ML. Our thorough analysis encompassed a substantial pool of 1,903 SCRA papers, ultimately narrowing our focus to a refined subset of 54 papers. This endeavor has yielded significant insights and key findings, making substantial contributions to the SCRA literature.
@@
    Our review thoroughly analyzes the existing state of research concerning the application of AI and ML techniques in SCRM. We have meticulously assessed a substantial body of literature to extract meaningful insights, specifically from 54 selected articles.
@@
    Through our analysis, we have identified the AI and ML techniques most commonly utilized in the domain of SCRA. This clarifies the methodologies employed and underscores the trends shaping the field.
@@
    Our review highlights the key findings and trends prevalent in the literature, shedding light on the progress, developments, and major areas of focus within AI and ML in SCRM.
@@
Despite these contributions, our research has some limitations. Our data primarily originated from Google Scholar and WoS rather than Scopus, potentially introducing biases. Our focus on papers published between 2015 and 2025 presents a limitation, as it may omit earlier relevant research that could provide valuable historical context. Additionally, our restriction to English-language papers may have omitted valuable non-English research.